# Prospects of constraining $f(T)$ gravity with the third–generation gravitational–wave detectors

### 03 Jan 2024 v1

@@ Abstract
Mergers of binary compact objects, accompanied with electromagnetic (EM) counterparts, offer excellent opportunities to explore varied cosmological models, since gravitational waves (GW) and EM counterparts always carry the information of luminosity distance and redshift, respectively. $f(T)$ gravity, which alters the background evolution and provides a friction term in the propagation of GW, can be tested by comparing the modified GW luminosity distance with the EM luminosity distance. Considering the third–generation gravitational–wave detectors, Einstein Telescope and two Cosmic Explorers, we simulate a series of GW events of binary neutron stars (BNS) and neutron-star-black-hole (NSBH) binary with EM counterparts. These simulations can be used to constrain $f(T)$ gravity (specially the Power-law model $f(T)=T+\alpha(-T)^{\beta}$ in this work) and other cosmological parameters, such as $\beta$ and Hubble constant. In addition, combining simulations with current observations of type Ia supernovae and baryon acoustic oscillations, we obtain tighter limitations for $f(T)$ gravity. We find that the estimated precision significantly improved when all three data sets are combined ($\Delta\beta\sim 0.03$), compared to analyzing the current observations alone ($\Delta\beta\sim 0.3$). Simultaneously, the uncertainty of the Hubble constant can be reduced to approximately $1\%$.
@@ I Introduction
General Relativity (GR) has been proven by many experiments, including the recent direct detection of Gravitational wave (GW) from the mergers of two compact objects [1, 2]. However, the presence of singularities at the centers of black holes and the breakdown of the equivalence principle at the singularities indicate that GR is not a universal theory of space-time [3], and the theoretical and observational challenges faced by the $\Lambda$CDM model might suggest a new theory of gravity [4, 5]. In addition to dark energy, many modified gravity models have been developed to account for the accelerating expansion of the universe. Among these models, one prominent approach involves incorporating the curvature through $f(R)$ gravity [6, 7]. As an extension of Teleparallel Equivalent of General Relativity (TEGR) [8], $f(T)$ gravity represented the most general modified form of TEGR and garnered considerable attention in previous studies [9, 10, 11]. Diverging from $f(R)$ gravity, characterized by curvature, $f(T)$ gravity utilizes torsion as a geometric object to describe gravity. Besides, $f(T)$ gravity is a second-order theory due to the fact that the torsion scalar $T$ depends only on the first derivatives of the vierbein [12]. Some studies have constrained $f(T)$ gravity by observations, such as the investigation with GW170817/GRB 170817A [13], the corresponding limitations based on cosmic microwave background (CMB) observations including the effects of primordial gravitational waves on the BB spectrum [14] and the impact of the linear scalar perturbations evolution on TT anisotropy power spectrum [15], and etc (see Ref. [16, 17, 18, 19] for comprehensive reviews).
@@
Within the framework of modified gravity and in the absence of anisotropic stress, the tensor perturbation bring the modifications of GW propagation equation [20]. In general, these modifications can alter the dispersion relationship, resulting in deviations between the speed of GW propagation and the speed of light. Recent studies of GW170817/GRB 170817A event [21, 22] provided a tighe constraint $\lvert c_{\text{gw}}/c-1\rvert<10^{-15}$ at low redshifts. However in $f(T)$ gravity, the sole modification in the equation of GW propagation is the friction term, which does not change the speed of GW propagation. Moreover, extensive investigations [23, 24, 25, 26, 27, 28, 29] devote to exploring the apparent difference between the GW luminosity distance ($d_{\rm L}^{gw}$) and electromagnetic (EM) luminosity distance ($d_{L}^{em}$), which result from the friction term in GW propagation function of modified gravity. Meanwhile, the effective field theory of $f(T)$ gravity within the cosmological framework presents an opportunity to modify the evolution of Hubble parameter $H(z)$, offering a model of late-time effective dark energy and releasing the $H_{0}$ tension [30, 31]. Besides, GW standard sirens present an independent and distinctive approach to investigating the $H_{0}$.
@@
In this paper, we present an end-to-end analysis of the simulated binary neutron stars (BNS) and neutron-star-black-hole (NSBH) binary samples with the third-generation gravitational-wave observatories, Einstein Telescope (ET) and Cosmic Explorer (CE). These prospective detectors are expected to exhibit significant improvements in sensitivity compared to current ground-based gravitational-wave detectors during the O5 observing period  [32, 33, 34], thereby providing a valuable opportunity to constrain $f(T)$ gravity at high redshifts. Because of the degeneracy between the inclination angle and the luminosity distance in GW observations, we construct different population models for BNS and NSBH events, including the presence of short gamma-ray bursts (SGRBs). Such EM counterpart improve the precisions of estimated $d_{\rm L}^{gw}$ [35], ultimately enhancing our ability to constrain $f(T)$ gravity from GW events. To establish the baselines of simulated GW, we begin by constraining various cosmological parameters using current observations, specifically Type Ia supernovae (SN Ia) and baryon acoustic oscillations (BAO). This initial analysis allows us to obtain reliable cosmological constraints. Then, we discuss the capability of constraining $f(T)$ gravity with the simulated GW catalog. By combining the information from SN Ia, BAO, and the simulated GW events, we find that the precision of constraining $f(T)$ gravity is significantly improved. This enhanced precision offers an opportunity to explore the possible deviations from GR within the framework of $f(T)$ gravity. It is worth noting that compared with previous works [16, 17], we have conducted a comprehensive analysis of simulated GW events, including the specific parameters of the binaries, and the conditions for generating the electromagnetic counterpart, etc. In addition, our simulations select a special class of GW events, which have small inclination angles and are accompanied by on-axis GRBs. In this case, the inclination angle can be limited to $\Delta\iota\leq 0.1$ rad directly. Therefore, the estimations of GW luminosity distance $\Delta d_{\rm L}$ will be more accurate, because the degeneracy between the inclination angle and the luminosity distance in GW analysis can be broken.
@@
The paper is organized as follows: In Sect. II, we provide a brief review of $f(T)$ gravity and discuss the modified propagation of GW within this framework. We focus on the Power-law model, which is one of the most commonly used models for $f(T)$. In Sect. III, we describe the baselines of GW simulations, which are based on realistic observations of SN Ia and BAO. In Sect. IV, we generate a simulated catalog of BNS and NSBH events as standard sirens, with detailed physical motivation and special selection criteria. Additionally, we analytically calculate the prospective accuracy of the luminosity distance estimation from GW using the combination of ET+2CE. In Sect. V, we present the main results and engage in in-depth discussions regarding the constraints of $f(T)$ gravity. Finally, we summarize the key findings of the study and present our conclusions in Sect. VI.
@@ II $f(T)$ gravity
In this section, we provide a brief overview of $f(T)$ gravity, along with an analytical treatment of the evolution of Hubble parameter and modified GW propagation within the framework of the Power-law model.
@@ II.1 Dynamical evolution
$\dot{H}$ =−$=-\frac{4\pi G\left(\rho_{m}+p_{m}\right)}{1+F_{T}+2TF_{TT}},$
@@
Once vierbein is obtained, the relation for the torsion scalar can be written as
$T$ =−$=-6H^{2}$ (4)
@@
where $T^{\rho}{}_{\mu\nu}=e_{A}^{\rho}\left(\partial_{\mu}e_{\nu}^{A}-\partial_{\nu}e_{\mu}^{A}\right)$ and $S_{\rho}{}^{\mu\nu}=\frac{1}{2}\left(\mathcal{K}^{\mu\nu}{}_{\rho}+\delta_{\rho}^{\mu}T^{\alpha\nu}{}_{\alpha}-\delta_{\rho}^{\nu}T^{\alpha\mu}{}_{\alpha}\right)$ represent the torsion tensor and the super-potential, respectively. The contortion tensor, denoted as $\mathcal{K}^{\mu\nu}{}_{\rho}=-\frac{1}{2}\left(T^{\mu\nu}{}_{\rho}-T^{\nu\mu}{}_{\rho}-T_{\rho}{}^{\mu\nu}\right)$, represents the divergence of the Levi-Civita connection in Riemannian spacetime and the Weitzenböck connection in Weitzenböck spacetime [8, 9].
@@ II.2 Power-law model
The evolution equation of $H(z)$, with the definition of normalized Hubble parameter denoted by $E^{2}(z)$, is expressed as [37]
@@
Here, we ignored the radiation density parameter of the current universe (i.e., $\Omega_{r}^{(0)}=0$). Fig. 1 showed the evolution of Eq. (6) compared with $\Lambda\mathrm{CDM}$ model ($\beta=0$). The second Friedmann equation can be written as
@@
where $\dot{H}$ denotes a derivative with respect to the cosmic time. Particularly, for $\Lambda\mathrm{CDM}$ model ($\beta=0$), Eq. (7) degenerates to
@@ II.3 GW propagation in $f(T)$ gravity
The perturbation part of the tensor sectors of $f(T)$ gravity around the FRW cosmological background leads to the equation of motion for the GW, which is consistent with the result from Arnowitt-Deser-Misner (ADM) decomposition of the vierbein field [38, 13, 18]
@@
where the prime denotes derivative with respect to the conformal time and $\lambda$ represents the two polarisation states of the tensor modes [39]. The dimensionless parameter $\beta_{T}$ is defined as
[…]
In Power-law model of $f(T)$ gravity, we introduce a simplifying factor $\kappa=\frac{(4-2\beta^{-1})}{3(\Omega_{m}^{(0)}-1)}$ and denote $\beta_{T}$ with $z$ as its variable. By substituting Eqs. (6) and (7) into Eq. (10), we obtain the expression of $\beta_{T}(z)$ as follows:
@@ III Current observations and GW-simulation baselines
$f(T)$ gravity provides a plausible explanation for cosmic evolution in various aspects by introducing “effective dark energy” during the late stages of the universe. It also changes structure formation in the early universe. This framework offers a promising avenue to alleviate the Hubble tension [30, 31]. In this section, we introduce the observations of SN Ia and BAO, which can be used to constrain the $f(T)$ gravity.
III.1 Type Ia Supernova
@@
where $H$ corresponds to the Hubble parameter in $f(T)$ gravity as defined in Equation Eq. (6). Additionally, the apparent magnitude of each SN Ia needs to be calibrated using an arbitrary fiducial absolute magnitude $M$. In our Markov Chain Monte Carlo (MCMC) analyses, we treat $M$ as a prior parameter and then marginalize it.
@@
∑$\sum_{ij}\left[m_{i}-m\left(z_{i}\right)\right]S_{\mathrm{sne},ij}^{-1}\left[m_{j}-m\left(z_{j}\right)\right],$
@@
Table 1: Injection parameters distributions of the BNS and NSBH populations
@@ III.2 Baryon Acoustic Oscillations
The measurements of baryon acoustic oscillations (BAO) have achieved an accuracy of $3\%$ across 14 narrow redshift shells. These measurements are based on a detailed analysis of galaxy data from the Baryon Oscillation Spectroscopic Survey (BOSS) Data Release 12 (DR12) and the extended Baryon Oscillation Spectroscopic Survey (eBOSS) Data Release 16 (DR16) in the range of redshift $0.660.32<z<$ 0.66 [49, 50, 51, 52, 53]. In our analysis, we primarily focus on the peak of the angular correlation function. The angular diameter distance is
@@
$\chi_{\mathrm{BAO}}^{2}\left(\Theta,r_{d}\right)$ =∑$=\sum_{ij}\left[\theta_{\mathrm{BAO}}\left(z_{i}\right)-\theta_{\mathrm{BAO},i}\right]$ (18)
@@ IV Modeling the gravitational signal
The merger rates of BNS and NSBH can be estimated from the GW Transient Catalog (GWTC) 1-3. The estimated BNS merger rate is $10-1700\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, while the NSBH merger rate is $7.8-140\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$. Previous studies by Jin et al.[54] found a higher BNS merger rate $452-2541~{}{\rm Gpc^{-3}~{}yr^{-1}}$, using sGRB data. In our work, we apply selection criteria based on the inclination angle, which results in approximately ∫$\int^{0.1}_{0}sin\iota{\rm d}\iota\sim\iota^{2}/2\sim 0.5\%$ events being selected from the total angular distribution. Considering the mean value of the merger rates (855 $\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ for BNS and 74 $\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ for NSBH), we expect there will be tens of thousands of BNS and NSBH events available over 1 year period, during the epoch with ET+2CE. Specifically, approximately 670 BNS and 50 NSBH events will have inclination angles $\iota\leq 0.1\,\rm rad$. Taking into account the proportion of black holes with high spin, which is approximately 24% [55], we estimate that there will be 1000 BNS (100 NSBH) events that can be observed within a two-year observation period. Due to the huge uncertainties associated with merger rates, we will subsequently engage in separate discussions concerning a set of 100 events and another set of 1000 events for BNS and NSBH, respectively. All of the injection parameters of BNS and NSBH simulations are summarized in Table 1. And, we formulate them in the following subsections, respectively.
Figure 3: The redshift distribution of the simulated 1000 BNS (left panel) and 1000 NSBH (right panel) events with ET+2CE, selected on the basis of different thresholds for the SNR.
IV.1 BNS
Since BNS are always accompanied with EM counterparts, we generate a population of simulated events in physically motivated parameter spaces. From this simulated population, we select a subset of events with small inclination angles to obtain well-estimated $d_{\rm L}^{gw}$.
∙$\bullet$ $\mathrm{NS}$ mass $M_{\mathrm{NS}}$-- The current understanding suggests that neutron stars in binary neutron star (BNS) systems have masses constrained within a narrow range, peaking at $1.33M_{\odot}$ [56, 57, 58]. Therefore, we assume that the distribution of NS mass in BNS systems follows an approximately Gaussian distribution, $M_{\mathrm{NS}}/M_{\odot}\sim\mathcal{N}\left(1.33,0.11^{2}\right)$.
∙$\bullet$ Equation of state (EoS) and Tidal deformability $\Lambda_{NS}$-- The AP4 EoS [59] is widely used in nuclear physics to describe the structure and properties of neutron stars. It incorporates various parameters and equations to account for the interactions between nucleons and other particles in dense matter. Relativistic effects and nuclear interactions are considered at high densities to accurately model the behavior of matter.

### 30 Jan 2024 v2

# Prospects of constraining $f(T)$ gravity with the third-generation gravitational-wave detectors

@@ Abstract
Mergers of binary compact objects, accompanied with electromagnetic (EM) counterparts, offer excellent opportunities to explore varied cosmological models, since gravitational waves (GWs) and EM counterparts always carry the information of luminosity distance and redshift, respectively. $f(T)$ gravity, which alters the background evolution and provides a friction term in the propagation of GWs, can be tested by comparing the modified GW luminosity distance with the EM luminosity distance. Considering the third-generation gravitational-wave detectors, Einstein Telescope and two cosmic explorers, we simulate a series of GW events of binary neutron stars and neutron-star–black-hole binaries with EM counterparts. These simulations can be used to constrain $f(T)$ gravity [especially the power-law model $f(T)=T+\alpha(-T)^{\beta}$ in this work] and other cosmological parameters, such as $\beta$ and the Hubble constant. In addition, combining simulations with current observations of type Ia supernovae and baryon acoustic oscillations, we obtain tighter limitations for $f(T)$ gravity. We find that the estimated precision significantly improved when all three datasets are combined ($\Delta\beta\sim 0.03$), compared to analyzing the current observations alone ($\Delta\beta\sim 0.3$). Simultaneously, the uncertainty of the Hubble constant can be reduced to approximately $1\%$.
@@ I Introduction
General relativity (GR) has been proven by many experiments, including the recent direct detection of gravitational waves (GWs) from the mergers of two compact objects [1, 2]. However, the presence of singularities at the centers of black holes and the breakdown of the equivalence principle at the singularities indicate that GR is not a universal theory of spacetime [3], and the theoretical and observational challenges faced by the $\Lambda$ cold dark matter ($\Lambda$CDM) model might suggest a new theory of gravity [4, 5]. In addition to dark energy, many modified gravity models have been developed to account for the accelerating expansion of the Universe. Among these models, one prominent approach involves incorporating the curvature through $f(R)$ gravity [6, 7]. As an extension of the teleparallel equivalent of general relativity (TEGR) [8], $f(T)$ gravity represented the most general modified form of TEGR and garnered considerable attention in previous studies [9, 10, 11]. Diverging from $f(R)$ gravity, characterized by curvature, $f(T)$ gravity utilizes torsion as a geometric object to describe gravity. In addition, $f(T)$ gravity is a second-order theory due to the fact that the torsion scalar $T$ depends only on the first derivatives of the vierbein [12]. Some studies have constrained $f(T)$ gravity by observations, such as the investigation with GW170817/$\gamma$-ray burst (GRB) 170817A [13], the corresponding limitations based on cosmic microwave background observations, including the effects of primordial gravitational waves on the BB spectrum [14], and the impact of the linear scalar perturbations evolution on the TT anisotropy power spectrum [15], etc. (see Refs. [16, 17, 18, 19] for comprehensive reviews).
@@
Within the framework of modified gravity and in the absence of anisotropic stress, the tensor perturbations bring the modifications of the GW propagation equation [20]. In general, these modifications can alter the dispersion relationship, resulting in deviations between the speed of GW propagation and the speed of light. Recent studies of the GW170817/GRB 170817A event [21, 22] provided a tight constraint $\lvert c_{\text{gw}}/c-1\rvert<10^{-15}$ at low redshifts. However in $f(T)$ gravity, the sole modification in the equation of GW propagation is the friction term, which does not change the speed of GW propagation. Moreover, extensive investigations [23, 24, 25, 26, 27, 28, 29] have been devoted to exploring the apparent difference between the GW luminosity distance ($d_{\rm L}^{gw}$) and electromagnetic (EM) luminosity distance ($d_{L}^{em}$), which result from the friction term in the GW propagation function of modified gravity. Meanwhile, the effective field theory of $f(T)$ gravity within the cosmological framework presents an opportunity to modify the evolution of Hubble parameter $H(z)$, offering a model of late-time effective dark energy and releasing the $H_{0}$ tension [30, 31]. Besides, GW standard sirens present an independent and distinctive approach to investigating the $H_{0}$.
@@
In this paper, we present an end-to-end analysis of the simulated binary neutron stars (BNSs) and neutron-star–black-hole (NSBH) binary samples with the third-generation gravitational-wave observatories, Einstein Telescope (ET) and Cosmic Explorer (CE). These prospective detectors are expected to exhibit significant improvements in sensitivity compared to current ground-based gravitational-wave detectors during the O5 observing period  [32, 33, 34], thereby providing a valuable opportunity to constrain $f(T)$ gravity at high redshifts. Because of the degeneracy between the inclination angle and the luminosity distance in GW observations, we construct different population models for BNS and NSBH events, including the presence of short $\gamma$-ray bursts (SGRBs). Such EM counterpart improves the precisions of estimated $d_{\rm L}^{gw}$ [35], ultimately enhancing our ability to constrain $f(T)$ gravity from GW events. To establish the baselines of simulated GWs, we begin by constraining various cosmological parameters using current observations, specifically type Ia supernovae (SN Ia) and baryon acoustic oscillations (BAOs). This initial analysis allows us to obtain reliable cosmological constraints. Then, we discuss the capability of constraining $f(T)$ gravity with the simulated GW catalog. By combining the information from SN Ia, BAOs, and the simulated GW events, we find that the precision of constraining $f(T)$ gravity is significantly improved. This enhanced precision offers an opportunity to explore the possible deviations from GR within the framework of $f(T)$ gravity. It is worth noting that, compared with previous works [16, 17], we have conducted a comprehensive analysis of simulated GW events, including the specific parameters of the binaries, and the conditions for generating the electromagnetic counterpart, etc. In addition, our simulations select a special class of GW events, which have small inclination angles and are accompanied by on axis GRBs. In this case, the inclination angle can be limited to $\Delta\iota\leq 0.1$ rad directly. Therefore, the estimations of GW luminosity distance $\Delta d_{\rm L}$ will be more accurate, because the degeneracy between the inclination angle and the luminosity distance in the GW analysis can be broken.
@@
The paper is organized as follows: In Sec. II, we provide a brief review of $f(T)$ gravity and discuss the modified propagation of GW within this framework. We focus on the power-law model, which is one of the most commonly used models for $f(T)$. In Sec. III, we describe the baselines of GW simulations, which are based on realistic observations of SN Ia and BAOs. In Sec. IV, we generate a simulated catalog of BNS and NSBH events as standard sirens, with detailed physical motivation and special selection criteria. Additionally, we analytically calculate the prospective accuracy of the luminosity distance estimation from GWs using the combination of ET+2CE. In Sec. V, we present the main results and engage in in-depth discussions regarding the constraints of $f(T)$ gravity. Finally, we summarize the key findings of the study and present our conclusions in Sec. VI.
@@ II $f(T)$ gravity
In this section, we provide a brief overview of $f(T)$ gravity, along with an analytical treatment of the evolution of the Hubble parameter and modified GW propagation within the framework of the power-law model.
@@ II.1 Dynamical evolution
$\dot{H}$ $=-\frac{4\pi G\left(\rho_{m}+p_{m}\right)}{1+F_{T}+2TF_{TT}},$
@@
Once the vierbein is obtained, the relation for the torsion scalar can be written as
$T$ $=-6H^{2}$ (4)
@@
where $T^{\rho}{}_{\mu\nu}=e_{A}^{\rho}\left(\partial_{\mu}e_{\nu}^{A}-\partial_{\nu}e_{\mu}^{A}\right)$ and $S_{\rho}{}^{\mu\nu}=\frac{1}{2}\left(\mathcal{K}^{\mu\nu}{}_{\rho}+\delta_{\rho}^{\mu}T^{\alpha\nu}{}_{\alpha}-\delta_{\rho}^{\nu}T^{\alpha\mu}{}_{\alpha}\right)$ represent the torsion tensor and the superpotential, respectively. The contortion tensor, denoted as $\mathcal{K}^{\mu\nu}{}_{\rho}=-\frac{1}{2}\left(T^{\mu\nu}{}_{\rho}-T^{\nu\mu}{}_{\rho}-T_{\rho}{}^{\mu\nu}\right)$, represents the divergence of the Levi-Civita connection in Riemannian spacetime and the Weitzenböck connection in Weitzenböck spacetime [8, 9].
@@ II.2 Power-law model
The evolution equation of $H(z)$, with the definition of the normalized Hubble parameter denoted by $E^{2}(z)$, is expressed as [37]
@@
Here, we ignored the radiation density parameter of the current Universe (i.e., $\Omega_{r}^{(0)}=0$). Figure. 1 showed the evolution of Eq. (6) compared with the $\Lambda\mathrm{CDM}$ model ($\beta=0$). The second Friedmann equation can be written as
@@
where $\dot{H}$ denotes a derivative with respect to the cosmic time. Particularly, for the $\Lambda\mathrm{CDM}$ model ($\beta=0$), Eq. (7) degenerates to
@@ II.3 GW propagation in $f(T)$ gravity
The perturbation part of the tensor sectors of $f(T)$ gravity around the FRW cosmological background leads to the equation of motion for the GW, which is consistent with the result from Arnowitt-Deser-Misner decomposition of the vierbein field [38, 13, 18]
$h_{\bm{\lambda}}^{\prime\prime}+2\mathcal{H}\left(1-\beta_{T}\right)h_{\bm{\lambda}}^{\prime}+k^{2}h_{\bm{\lambda}}=0,$ (9)
where the prime denotes derivative with respect to the conformal time and $\lambda$ represents the two polarization states of the tensor modes [39]. The dimensionless parameter $\beta_{T}$ is defined as
$\beta_{T}=-\frac{f_{T}^{\prime}}{2\mathcal{H}f_{T}}=-\frac{\dot{f_{T}}}{2Hf_{T}}=\frac{6f_{TT}\dot{H}}{f_{T}}.$ (10)
By resubstituting the scale factor $\frac{\tilde{a}^{\prime}}{\tilde{a}}=\mathcal{H}[1-\beta_{T}]$ and performing some straightforward derivations, the redefined GW luminosity distance is related to the electromagnetic luminosity distance  [40]
@@
In the power-law model of $f(T)$ gravity, we introduce a simplifying factor $\kappa=\frac{(4-2\beta^{-1})}{3(\Omega_{m}^{(0)}-1)}$ and denote $\beta_{T}$ with $z$ as its variable. By substituting Eqs. (6) and (7) into Eq. (10), we obtain the expression of $\beta_{T}(z)$ as follows:
@@ III Current observations and GW-simulation baselines
$f(T)$ gravity provides a plausible explanation for cosmic evolution in various aspects by introducing effective dark energy during the late stages of the Universe. It also changes structure formation in the early Universe. This framework offers a promising avenue to alleviate the Hubble tension [30, 31]. In this section, we introduce the observations of SN Ia and BAOs, which can be used to constrain the $f(T)$ gravity.
III.1 Type Ia supernovae
@@
where $H$ corresponds to the Hubble parameter in $f(T)$ gravity as defined in Eq. (6). Additionally, the apparent magnitude of each SN Ia needs to be calibrated using an arbitrary fiducial absolute magnitude $M$. In our Markov chain Monte Carlo analyses, we treat $M$ as a prior parameter and then marginalize it.
@@
$\sum_{ij}\left[m_{i}-m\left(z_{i}\right)\right]S_{\mathrm{sne},ij}^{-1}\left[m_{j}-m\left(z_{j}\right)\right],$
@@
Table 1: Injection parameter distributions of the BNS and NSBH populations
@@ III.2 Baryon acoustic oscillations
The measurements of BAOs have achieved an accuracy of $3\%$ across 14 narrow redshift shells. These measurements are based on a detailed analysis of Galaxy data from BOSS data release 12 and eBOSS data release 16 in the range of redshift $0.660.32<z<$ 0.66 [49, 50, 51, 52, 53]. In our analysis, we primarily focus on the peak of the angular correlation function. The angular diameter distance is
@@
$\chi_{\mathrm{BAO}}^{2}\left(\Theta,r_{d}\right)$ $=\sum_{ij}\left[\theta_{\mathrm{BAO}}\left(z_{i}\right)-\theta_{\mathrm{BAO},i}\right]$ (18)
@@ IV Modeling the gravitational signal
The merger rates of BNSs and NSBH can be estimated from the GW transient catalog 1–3. The estimated BNS merger rate is 10170010170010–170010 – 1700 $\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, while the NSBH merger rate is 7.81407.81407.8–1407.8 – 140 $\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$. Previous studies by Jin et al.[54] found a higher BNS merger rate 45225414522541452–2541452 – 2541 $\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, using SGRB data. In our work, we apply selection criteria based on the inclination angle, which results in approximately $\int^{0.1}_{0}\sin\iota{\rm d}\iota\sim\iota^{2}/2\sim 0.5\%$ events being selected from the total angular distribution. Considering the mean value of the merger rates (855 $\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ for BNSs and 74 $\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ for NSBH), we expect there will be tens of thousands of BNS and NSBH events available over a one year period, during the epoch with ET+2CE. Specifically, approximately 670 BNS and 50 NSBH events will have inclination angles $\iota\leq 0.1\,\rm rad$. Taking into account the proportion of black holes with high spin, which is approximately 24% [55], we estimate that there will be 1000 BNS (100 NSBH) events that can be observed within a two year observation period. Because of the huge uncertainties associated with merger rates, we will subsequently engage in separate discussions concerning a set of 100 events and another set of 1000 events for BNSs and NSBH, respectively. All of the injection parameters of BNS and NSBH simulations are summarized in Table 1, and we formulate them in the following subsections, respectively.
Figure 3: The redshift distribution of the simulated 1000 BNS (left) and 1000 NSBH (right) events with ET+2CE, selected on the basis of different thresholds for the SNR.
IV.1 BNSs
Since BNSs are always accompanied with EM counterparts, we generate a population of simulated events in physically motivated parameter spaces. From this simulated population, we select a subset of events with small inclination angles to obtain well-estimated $d_{\rm L}^{gw}$.
(i) Neutron star (NS) mass $M_{\mathrm{NS}}$: The current understanding suggests that neutron stars in BNS systems have masses constrained within a narrow range, peaking at $1.33\mathrm{M}_{\odot}$ [56, 57, 58]. Therefore, we assume that the distribution of NS mass in BNS systems follows an approximately Gaussian distribution, $M_{\mathrm{NS}}/\mathrm{M}_{\odot}\sim\mathcal{N}\left(1.33,0.11^{2}\right)$.
(ii) Equation of state (EOS) and tidal deformability $\Lambda_{NS}$: The AP4 EOS [59] is widely used in nuclear physics to describe the structure and properties of neutron stars. It incorporates various parameters and equations to account for the interactions between nucleons and other particles in dense matter. Relativistic effects and nuclear interactions are considered at high densities to accurately model the behavior of matter.